how do i solve this: "Suppose that $5 is invested at 4%, compounded quarterly. How much would be in the account after 300 years? Give your answer to the nearest cent.
Answer by stanbon(75887)

Suppose that $5 is invested at 4%, compounded quarterly. How much would be in the account after 300 years?
---
A(t) = P(1+(r/n))^(nt)
----
A(300) = 5(1+(0.04/4))^(4*300)
----
A(300) = 5(1.01)^1200
---
A(300) = $766,687.78
========================
Cheers,
Stan H.
===============